# Spatial-Sign based Maxsum Test for High Dimensional Location Parameters

### v1

@@ 1 Introduction
In practical scenarios, it’s often unknown whether the alternative is dense or sparse. This has led to numerous studies proposing an adaptive strategy that combines the sum-type test and max-type test. For high-dimensional mean problems, Xu et al. (2016) integrated different $L_{r}$-norms of the sample means. He et al. (2021) introduced a family of U-statistics as an unbiased estimator of the $L_{r}$-norm of the mean vectors, demonstrating that U-statistics of different finite orders are asymptotically independent, normally distributed, and also independent from the maximum-type test statistic. Feng et al. (2022a) relaxed the covariance matrix assumption to establish independence between the sum-type test statistic and the max-type test statistic. There are also many other studies showing the asymptotic independence between the sum-type test statistics and the max-type test statistic for other high-dimensional problems. For instance, Wu et al. (2019) and Wu et al. (2020) examined the coefficient test in high-dimensional generalized linear models. Feng et al. (2022b) looked at the cross-sectional independence test in high-dimensional panel data models. Yu et al. (2022) focused on testing the high-dimensional covariance matrix. Feng et al. (2022d) considered the high-dimensional white noise test, while Wang and Feng (2023) looked at high-dimensional change point inference. Ma et al. (2023) and Yu et al. (2023) considered testing the alpha of high-dimensional time-varying and linear factor pricing models, respectively.
@@ 2.1 Bahadur representation and Gaussian approximation
$\frac{1}{n}\sum_{i=1}^{n}U\left(\bm{\epsilon}_{i}\right)=0\text{ and }\frac{p}{n}\operatorname{diag}\left\{\sum_{i=1}^{n}U\left(\bm{\epsilon}_{i}\right)U\left(\bm{\epsilon}_{i}\right)^{T}\right\}=\mathbf{I}_{p},$ (2.4)
where $\bm{\epsilon}_{i}=\mathbf{D}^{-1/2}\left(\bm{X}_{i}-\bm{\theta}\right)$. ($\mathbf{D}$, $\left.\bm{\theta}\right)$ can be viewed as a simplified version of HettmanspergerRandles (HR) estimator without considering the off-diagonal elements of $\mathbf{S}$. We can adapt the recursive algorithm of Feng et al. (2016) to solve Equation 2.4. That is, repeat the following three steps until convergence:
(i) $\bm{\epsilon}_{i}\leftarrow\mathbf{D}^{-1/2}\left(\mathbf{X}_{i}-\bm{\theta}\right),\quad j=1,\cdots,n_{i}$;
@@
(iii) $\mathbf{D}\leftarrow p\mathbf{D}^{1/2}\operatorname{diag}\left\{n^{-1}\sum_{j=1}^{n}U\left(\bm{\epsilon}_{i}\right)U\left(\bm{\epsilon}_{i}\right)^{T}\right\}\mathbf{D}^{1/2}$.
The resulting estimators of location and diagonal matrix are denoted as $\hat{\bm{\theta}}$ and $\hat{\mathbf{D}}$. The sample mean and sample variances can be utilized as initial estimators. Regrettably, no evidence has been found to confirm the convergence of the aforementioned algorithm, even in low-dimensional scenarios, despite its consistent practical effectiveness. The existence or uniqueness of the HR estimate mentioned above also lacks proof. This topic certainly warrants further investigation.
In this section, we investigate some theoretical properties based on maximum-norm about $\hat{\bm{\theta}}$. Similar with the proof of Lemma 1 and Theorem 1 in Cheng et al. (2023), we give the Bahadur representation of $\hat{\bf D}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})$ and the Gaussian approximation result for $\hat{\bf D}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})$ over hyperrectangles. Based on Gaussian approximation, we can easily derive the limiting distribution of $\hat{\bm{\theta}}$ based on the maximum-norm.
@@
The shape matrix ${\bf R}=\mathbf{D}^{-1/2}\Gamma\Gamma^{\top}\mathbf{D}^{-1/2}=\left(\sigma_{j\ell}\right)_{p\times p}$ satisfies $\max_{j=1,\cdots,p}\sum_{\ell=1}^{p}\left|\sigma_{j\ell}\right|\leqslant a_{0}(p)$. In addition, $\lim\inf_{p\rightarrow\infty}\min_{j=1,2,\cdots,p}d_{j}>\underline{d}$ for some constant $\underline{d}>0$, where $\mathbf{D}=\operatorname{diag}\{d_{1}^{2},d_{2}^{2},\cdots,d_{p}^{2}\}$.
Remark 1  Assumption 1 is the same as Condition C.1 in Cheng et al. (2023), which ensure that $\bm{\theta}$ in model (2.1) is the population spatial median and $W_{i,j}$ has a sub-exponential distribution. If $\bm{W}_{i}\sim N(\bm{0},{\bf I}_{p})$, $\bm{X}_{i}$ follows a elliptical symmetric distribution. Assumption 2 extend the Assumption 1 in Zou et al. (2014), which indicates that $\zeta_{k}\asymp p^{-k/2}$. Assumption 3 means the correlation between those variables could not be too large, which is similar to the matrix class in Bickel and Levina (2008).
@@
(Bahadur representation) Under Assumption 1-3 with $a_{0}(p)\asymp p^{1-\delta}$, if $\log p=o(n^{1/3})$ and $\log n=o(p^{1/3\wedge\delta})$, we have that:
@@
$\|C_{n}\|_{\infty}=O_{p}(n^{-1/4}\log^{1/2}(np)+p^{-(1/6\wedge\delta/2)}\log^{1/2}(np)+n^{-1/2}(\log p)^{1/2}\log^{1/2}(np)).$
@@
Let $\mathcal{A}^{\text{re }}=\left\{\prod_{j=1}^{p}\left[a_{j},b_{j}\right]:-\infty\leqslant a_{j}\leqslant b_{j}\leqslant\infty,j=1,\ldots,p\right\}$ be the class of rectangles in $\mathbb{R}^{p}$. Based on the Bahadur representation of $\hat{\bm{\theta}}_{n}$, we acquire the following Gaussian approximation of $\hat{\mathbf{D}}^{-1/2}\left(\hat{\bm{\theta}}-\bm{\theta}\right)$ in rectangle $\mathcal{A}^{\text{re }}$.
@@
(Gaussian approximation) Assume Assumption 1-3 with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leqslant 1/2$ hold. If $\log p=o\left(n^{1/5}\right)$ and $\log n=o\left(p^{1/3\wedge\delta}\right)$, then
@@
Suppose the conditions assumed in Lemma 2 hold, as $n\rightarrow\infty$,
@@
(Variance approximation) Suppose $\bm{G}\sim N\left(0,\zeta_{1}^{-2}\mathbf{\Sigma}_{u}\right)$ and $\bm{Z}\sim N\left(0,\zeta_{1}^{-2}p^{-1}\mathbf{R}\right)$, as $(n,p)\rightarrow\infty$,
@@ 2.2 Max-type test procedure
Let ${\bf R}=\left(\sigma_{ij}\right)_{1\leq i,j\leq p}$. For some $\varrho\in(0,1)$, assume $|\sigma_{ij}|\leq\varrho$ for all $1\leq i<j\leq p$ and $p\geq 2$. Suppose $\left\{\delta_{p};p\geq 1\right\}$ and $\left\{\kappa_{p};p\geq 1\right\}$ are positive constants with $\delta_{p}=o(1/\log p)$ and $\kappa=\kappa_{p}\rightarrow 0$ as $p\rightarrow\infty$. For $1\leq i\leq p$, define $B_{p,i}=\left\{1\leq j\leq p;\left|\sigma_{ij}\right|\geq\delta_{p}\right\}$ and $C_{p}=\left\{1\leq i\leq p;\left|B_{p,i}\right|\geq p^{\kappa}\right\}$. We assume that $\left|C_{p}\right|/p\rightarrow 0$ as $p\rightarrow\infty$.
@@
Suppose Assumption 1-4 holds, by the Theorem 2 in Feng et al. (2022a), we can see that $p\zeta_{1}^{2}\max_{1\leq i\leq p}Z_{i}^{2}-2\log p+\log\log p$ converges to a Gumbel distribution with cdf $F(x)=\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-x/2}\right\}$ as $p\rightarrow\infty$. In combining with the Theorem 1 we can conclude that,
@@
Next we replace $E\left(R^{-1}\right)$ with its estimators. Denote $\hat{R}_{i}=\left\|\hat{\mathbf{D}}^{-1/2}\left(\mathbf{X}_{i}-\hat{\bm{\theta}}\right)\right\|$. Then the estimator is defined as $\hat{E\left(R_{i}^{-1}\right)}=\frac{1}{n}\sum_{i=1}^{n}\hat{R}_{i}^{-1}:=\hat{\zeta}_{1}$, and the proof of consistency is shown in Lemma 7. Because the convergence rate of maximum is very slow, we propose a adjust max-type test statistic which based on the scalar-invariant spatial median $\hat{\bm{\theta}}$,
@@
Suppose the conditions in Theorem 1 and Assumption 4 hold. Under the null hypothesis, as $(n,p)\to\infty$, we have
@@
Suppose the conditions assumed in Theorem 2 holds. For any given $\alpha\in(0,1)$, if $\|\bm{\theta}\|_{\infty}\geq\widetilde{C}n^{-1/2}\{\log p-2\log\log(1-\alpha)^{-1}\}^{1/2}$, for some large enough constant $\widetilde{C}$, then
@@
Given a fixed significance level $\alpha$, the test $T_{MAX}$ attains consistency if $\|\bm{\theta}\|_{\infty}\geq\widetilde{C}\sqrt{\log p/n}$, provided that $\widetilde{C}$ is sufficiently large. This is the minimax rate optimal for testing against sparse alternatives, as stated in Theorem 3 of Cai et al. (2013). If $\widetilde{C}$ is adequately small, then it becomes impossible for any $\alpha$-level test to reject the null hypothesis with a probability approaching one. Therefore, Theorem 3 also demonstrates the optimality of our proposed test $T_{MAX}$.
@@
$H_{1}:\bm{\theta}=(\theta_{1},0,\cdots,0)^{T},\theta_{1}>0$
which means there are only one variable has nonzero means. Let $x_{\alpha}=2\log p-\log\log p+q_{1-\alpha}$. In this case,
@@ 3.1 Exisiting Sum-type test
$T_{SUM}=\frac{2}{n(n-1)}{\sum\sum}_{i<j}U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\bm{X}_{i}\right)^{T}U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\bm{X}_{j}\right)$
@@
Under Assumptions 1-3,6-7. and $H_{0}$, as $\min\left(p,n\right)\rightarrow\infty,T_{SUM}/\sigma_{n}\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1)$.
@@
$\widehat{\operatorname{tr}\left(\mathbf{R}^{2}\right)}=\frac{p^{2}}{n(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}^{n}\left(U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\left(\bm{X}_{i}-\hat{\bm{\theta}}_{ij}\right)\right)^{T}U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\left(\bm{X}_{j}-\hat{\bm{\theta}}_{ij}\right)\right)\right)^{2},$
@@
By Theorem 4 and 5, the asymptotic power function of $T_{SUM}$ is

### v2

@@ 1 Introduction
In practical scenarios, it’s often unknown whether the alternative is dense or sparse. This has led to numerous studies proposing an adaptive strategy that combines the sum-type test and max-type test. For high-dimensional mean problems, Xu et al. (2016) integrated different $L_{r}$-norms of the sample means. He et al. (2021) introduced a family of U-statistics as an unbiased estimator of the $L_{r}$-norm of the mean vectors, covariance matrices and regression coefficients, demonstrating that U-statistics of different finite orders are asymptotically independent, normally distributed, and also independent from the maximum-type test statistic. Feng et al. (2022a) relaxed the covariance matrix assumption to establish independence between the sum-type test statistic and the max-type test statistic. There are also many other studies showing the asymptotic independence between the sum-type test statistics and the max-type test statistic for other high-dimensional problems. For instance, Wu et al. (2019) and Wu et al. (2020) examined the coefficient test in high-dimensional generalized linear models. Feng et al. (2022b) looked at the cross-sectional independence test in high-dimensional panel data models. Yu et al. (2022) focused on testing the high-dimensional covariance matrix. Feng et al. (2022d) considered the high-dimensional white noise test, while Wang and Feng (2023) looked at high-dimensional change point inference. Ma et al. (2023) and Yu et al. (2023) considered testing the alpha of high-dimensional time-varying and linear factor pricing models, respectively.
@@ 2.1 Bahadur representation and Gaussian approximation
$\frac{1}{n}\sum_{i=1}^{n}U\left(\bm{\epsilon}_{i}\right)=0\text{ and }\frac{p}{n}\operatorname{diag}\left\{\sum_{i=1}^{n}U\left(\bm{\epsilon}_{i}\right)U\left(\bm{\epsilon}_{i}\right)^{\top}\right\}=\mathbf{I}_{p},$ (2.4)
where $\bm{\epsilon}_{i}=\mathbf{D}^{-1/2}\left(\bm{X}_{i}-\bm{\theta}\right)$. ($\mathbf{D}$, $\left.\bm{\theta}\right)$ can be viewed as a simplified version of HettmanspergerRandles (HR) estimator , which is proposed in Hettmansperger and Randles (2002), without considering the off-diagonal elements of $\mathbf{S}$. We can adapt the recursive algorithm of Feng et al. (2016) to solve Equation 2.4. That is, repeat the following three steps until convergence:
(i) $\bm{\epsilon}_{i}\leftarrow\mathbf{D}^{-1/2}\left(\mathbf{X}_{i}-\bm{\theta}\right),\quad j=1,\cdots,n$;
@@
(iii) $\mathbf{D}\leftarrow p\mathbf{D}^{1/2}\operatorname{diag}\left\{n^{-1}\sum_{j=1}^{n}U\left(\bm{\epsilon}_{i}\right)U\left(\bm{\epsilon}_{i}\right)^{\top}\right\}\mathbf{D}^{1/2}$.
The resulting estimators of location and diagonal matrix are denoted as $\hat{\bm{\theta}}$ and $\hat{\mathbf{D}}$. The sample mean and sample variances can be utilized as initial estimators. Regrettably, no evidence has been found to confirm the convergence of the aforementioned algorithm, even in low-dimensional scenarios, despite its consistent practical effectiveness. The existence or uniqueness of the HR estimator mentioned above also lacks proof. This topic certainly warrants further investigation.
In this section, we investigate some theoretical properties based on maximum-norm about $\hat{\bm{\theta}}$. Similar to the proof of Lemma 1 and Theorem 1 in Cheng et al. (2023), we give the Bahadur representation of $\hat{\bf D}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})$ and the Gaussian approximation result for $\hat{\bf D}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})$ over hyperrectangles. Based on Gaussian approximation, we can easily derive the limiting distribution of $\hat{\bm{\theta}}$ based on the maximum-norm.
@@
The shape matrix ${\bf R}=\mathbf{D}^{-1/2}\mathbf{\Gamma}\mathbf{\Gamma}^{\top}\mathbf{D}^{-1/2}=\left(\sigma_{j\ell}\right)_{p\times p}$ satisfies $\max_{j=1,\cdots,p}\sum_{\ell=1}^{p}\left|\sigma_{j\ell}\right|\leqslant a_{0}(p)$ where $a_{0}(p)$ is a constant depending only on dimension $p$. In addition, $\lim\inf_{p\rightarrow\infty}\min_{j=1,2,\cdots,p}d_{j}>\underline{d}$ for some constant $\underline{d}>0$, where $\mathbf{D}=\operatorname{diag}\{d_{1}^{2},d_{2}^{2},\cdots,d_{p}^{2}\}$.
[…]
@@
(Bahadur representation) Assume Assumptions 1-3 with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leq 1/2$ hold. If $\log p=o(n^{1/3})$ and $\log n=o(p^{1/3\wedge\delta})$, then
@@
$\|C_{n}\|_{\infty}=O_{p}\{n^{-1/4}\log^{1/2}(np)+p^{-(1/6\wedge\delta/2)}\log^{1/2}(np)+n^{-1/2}(\log p)^{1/2}\log^{1/2}(np)\}.$
@@
Let $\mathcal{A}^{\text{re }}=\left\{\prod_{j=1}^{p}\left[a_{j},b_{j}\right]:-\infty\leqslant a_{j}\leqslant b_{j}\leqslant\infty,j=1,\ldots,p\right\}$ be the class of rectangles in $\mathbb{R}^{p}$. Based on the Bahadur representation of $\hat{\bm{\theta}}$, we acquire the following Gaussian approximation of $\hat{\mathbf{D}}^{-1/2}(\hat{\bm{\theta}}-\bm{\theta})$ in rectangle $\mathcal{A}^{\text{re }}$.
@@
(Gaussian approximation) Assume Assumptions 1-3 with $a_{0}(p)\asymp p^{1-\delta}$ for some positive constant $\delta\leqslant 1/2$ hold. If $\log p=o\left(n^{1/5}\right)$ and $\log n=o\left(p^{1/3\wedge\delta}\right)$, then
@@
Under the assumptions of Lemma 2, as $n\rightarrow\infty$, we have
@@
(Variance approximation) Suppose $\bm{G}\sim N\left(0,\zeta_{1}^{-2}\mathbf{\Sigma}_{u}\right)$ and $\bm{Z}\sim N\left(0,\zeta_{1}^{-2}p^{-1}\mathbf{R}\right)$, under the assumptions of Lemma 2, as $(n,p)\rightarrow\infty$, we have
@@ 2.2 Max-type test procedure
For some $\varrho\in(0,1)$, assume $|\sigma_{ij}|\leq\varrho$ for all $1\leq i<j\leq p$ and $p\geq 2$. Suppose $\left\{\delta_{p};p\geq 1\right\}$ and $\left\{\kappa_{p};p\geq 1\right\}$ are positive constants with $\delta_{p}=o(1/\log p)$ and $\kappa=\kappa_{p}\rightarrow 0$ as $p\rightarrow\infty$. For $1\leq i\leq p$, define $B_{p,i}=\left\{1\leq j\leq p;\left|\sigma_{ij}\right|\geq\delta_{p}\right\}$ and $C_{p}=\left\{1\leq i\leq p;\left|B_{p,i}\right|\geq p^{\kappa}\right\}$. We assume that $\left|C_{p}\right|/p\rightarrow 0$ as $p\rightarrow\infty$.
@@
Suppose Assumption 1-4 hold, by the Theorem 2 in Feng et al. (2022a), we can see that $p\zeta_{1}^{2}\max_{1\leq i\leq p}Z_{i}^{2}-2\log p+\log\log p$ converges to a Gumbel distribution with cdf $F(x)=\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-x/2}\right\}$ as $p\rightarrow\infty$. In combining with the Theorem 1 we can conclude that,
@@
Next we replace $E\left(R^{-1}\right)$ with its estimators. We denote $\hat{R}_{i}=\|\hat{\mathbf{D}}^{-1/2}(\mathbf{X}_{i}-\hat{\bm{\theta}})\|$. Then the estimator is defined as $\hat{\zeta}_{1}:=\frac{1}{n}\sum_{i=1}^{n}\hat{R}_{i}^{-1}$, and the proof of consistency is shown in Lemma 7. Because the convergence rate of maximum is very slow, we propose a adjust max-type test statistic which based on the scalar-invariant spatial median $\hat{\bm{\theta}}$,
@@
Suppose the assumptions in Theorem 1 and Assumption 4 hold. Under the null hypothesis, as $(n,p)\to\infty$, we have
@@
Suppose the Assumptions in Theorem 2 hold. For any given $\alpha\in(0,1)$, if $\|\bm{\theta}\|_{\infty}\geq\widetilde{C}n^{-1/2}\{\log p-2\log\log(1-\alpha)^{-1}\}^{1/2}$ for some large enough constant $\widetilde{C}$, we have
@@
Given a fixed significant level $\alpha$, the test $T_{MAX}$ attains consistency if $\|\bm{\theta}\|_{\infty}\geq\widetilde{C}\sqrt{\log p/n}$, provided that $\widetilde{C}$ is sufficiently large. This is the minimax rate optimal for testing against sparse alternatives, as stated in Theorem 3 of Cai et al. (2013). If $\widetilde{C}$ is adequately small, then it becomes impossible for any $\alpha$-level test to reject the null hypothesis with a probability approaching one. Therefore, Theorem 3 also demonstrates the optimality of our proposed test $T_{MAX}$.
@@
$H_{1}:\bm{\theta}=(\theta_{1},0,\cdots,0)^{\top},\theta_{1}>0,$
which means there are only one variable has nonzero mean. Let $x_{\alpha}=2\log p-\log\log p+q_{1-\alpha}$. In this case,
@@ 3.1 Exisiting Sum-type test
$T_{SUM}=\frac{2}{n(n-1)}{\sum\sum}_{i<j}U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\bm{X}_{i}\right)^{T}U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\bm{X}_{j}\right),$
@@
Remark 4  Assumption 6 is a common condition for sum-type test statistic in high dimensions, see Chen and Qin (2010); Feng et al. (2016); Feng and Sun (2016), which requires that the eigenvalues of ${\bf R}$ not diverge excessively. If all the eigenvalues of ${\bf R}$ are bounded, ${\rm tr}({\bf R}^{2})=O(p),{\rm tr}({\bf R}^{4})=O(p)$. So the Assumption 6 holds trivially. In this case, Assumption 7 becomes $p=O(n^{2})$ and $p/n\rightarrow\infty$. Actually, Assumption 7 indicates $p=O(n^{2})$ and it is not necessary for the eigenvalues to be bounded. For $b$ unbounded eigenvalues with respect dimension $p$, the sufficient condition for Assumption 6 is $\lambda_{(p)}/\lambda_{(1)}=o((p-b)^{1/2}b^{-1/4})$. For Assumption 7, a sufficient condition is $\lambda_{i}=1+o(p^{1/2}n^{-1/2})$. When the assumptions about the spectrum of ${\bf R}$ do not hold, there would be a bias therm in sum-type statistic that difficult to calculate and deserves to be investigated further , see the Supplemental Material of Feng et al. (2016) for more details.
@@
Under Assumptions 1-3,6-7 and $H_{0}$, as $\min\left(p,n\right)\rightarrow\infty,T_{SUM}/\sigma_{n}\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1)$.
@@
$\widehat{\operatorname{tr}\left({{\bf R}}^{2}\right)}=\frac{p^{2}}{n(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}^{n}\left(U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\left(\bm{X}_{i}-\hat{\bm{\theta}}_{ij}\right)\right)^{T}U\left(\hat{\mathbf{D}}_{ij}^{-1/2}\left(\bm{X}_{j}-\hat{\bm{\theta}}_{ij}\right)\right)\right)^{2},$
@@
By Theorem 4 and 5, the power function of $T_{SUM}$ can be approximated as